OpenStudy (anonymous):

Find x from the equation: 4^x + 2^x = 12. I GIVE MEDALS. PLEASE HELP ME :D

jimthompson5910 (jim_thompson5910):

let z = 2^x square both sides to get z^2 = (2^x)^2 z^2 = 2^(2x) z^2 = (2^2)^x z^2 = 4^x

jimthompson5910 (jim_thompson5910):

So because z = 2^x and z^2 = 4^x, we go from 4^x + 2^x = 12 to z^2 + z = 12

jimthompson5910 (jim_thompson5910):

Your job from here is to solve z^2 + z = 12 for z You'll get two solutions for z. Once you get those two solutions, use each one to find the solutions in terms of x.

jimthompson5910 (jim_thompson5910):

Tell me what you get when you solve z^2 + z = 12 for z

OpenStudy (anonymous):

@jim_thompson5910 i got z = 3 and z = -4 :)

OpenStudy (anonymous):

@jim_thompson5910 so, i will try, 3 = 2^x and -4 = 2^x right??

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

if 3 = 2^x, then x = ???

OpenStudy (anonymous):

@jim_thompson5910 3 = 2^x.. can it be solved without a scientific calculator or anything like logarithm or something?? :)

jimthompson5910 (jim_thompson5910):

it can be solved using logs

OpenStudy (anonymous):

@jim_thompson5910 from 3 = 2^x..

OpenStudy (anonymous):

it is log 3 to base 2

jimthompson5910 (jim_thompson5910):

if b^x = y, then log(b,y) = x where "log(b,y)" means "log base b of y"

jimthompson5910 (jim_thompson5910):

so 2^x = 3 turns into x = log(2,3)

OpenStudy (anonymous):

x = log 3 base 2 @jim_thompson5910 ? :)

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

@jim_thompson5910 ok, i will wait ;)

jimthompson5910 (jim_thompson5910):

check out the attached image

OpenStudy (anonymous):

@jim_thompson5910 yah yah i got the log thingy :)

jimthompson5910 (jim_thompson5910):

The other equation 2^x = -4 does not have a solution because 2^x is always positive (for any value of x)

OpenStudy (anonymous):

@jim_thompson5910 ok, got it. so the value of x is log(2,3)! :)) i tried it in sci calculator, the result is exactly 12 :)

jimthompson5910 (jim_thompson5910):

I'm not sure how you got log(2,3) to equal 12 when that's completely false

jimthompson5910 (jim_thompson5910):

But I think you meant to say that you plugged in x = log(2,3) into 4^x + 2^x = 12 and you got 12 = 12 when you reduced

OpenStudy (anonymous):

@jim_thompson5910 no no, what is mean is log 3 to base 2 like from your attachment

OpenStudy (anonymous):

2 is the base, i mean

jimthompson5910 (jim_thompson5910):

I gotcha

OpenStudy (anonymous):

@jim_thompson5910 thanks!! the 'equation' button isnt working either :(

jimthompson5910 (jim_thompson5910):

yeah which is why I had to attach the image